OpenStudy (anonymous):

Find the system determinant of the given matrix. -30 -6 6 30

OpenStudy (unklerhaukus):

-30, -6 6 , 30 ?

OpenStudy (unklerhaukus):

\[\mathbf M=\begin{bmatrix}a&b\\c&d\end{bmatrix}\] determinant of matrix M \[\det(\mathbf M)=\begin{vmatrix}a&b\\c&d\end{vmatrix}=ad-bc\]